Hemoglobin (C738H1166N812O203Fe) is a substance found in the red blood cell and is responsible for the absorption of Oxygen gas into our bloodstream by binding oxygen to the iron. Based on the chemical composition of the hemoglobin, how many percent of the substance is made up of iron? 

1
Expert's answer
2021-02-15T03:26:19-0500

Molar mass of Haemoglobin= (12x738) + (1x1166) + (14x812) + (16x203) + 56= 24694


Percentage by mass of Fe= 56/24694 x 100% = 0.23%